Circular No. 8894

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2007 V2 (HILL) R. E. Hill reports his discovery of a comet (discovery observation tabu-

lated below) of mag ∼ 19 with a diffuse 4′′ × 6′′ coma and a broad tail ∼ 2′ long in p.a. 260^{o} in four co-added 30-s CCD exposures taken in 2′′–3′′ seeing with the Mt. Lemmon 1.5-m reflector. His four co-added 60-s unfiltered images taken the following night (Nov. 10.3 UT) show a well-condensed cirular coma and a broad tail 2′.6 long. Following posting on the Minor Planet Center’s NEOCP webpage, J. Young reports that CCD images taken by W. Young and himself in very poor seeing on Nov. 10.24–10.28 with the Table Mountain 0.61-m f/16 Cassegrain reflector show a round 6′′ diffuse coma with with no central condensation and no visible tail seen. R. Ligustri and D. Da Rio (Talmassons, Udine, Italy, 0.35-m reflector) write that their twenty co-added 180-s exposures from Nov. 9.83–9.85 show a very faint coma of diameter ∼ 20′′.

2007 UT  α_{2000}  δ_{2000}
Nov. 9.29539  2^{h}38^{m}15.^{s}99  +12^{o}05′51′′.8

The available astrometry (with prediscovery Spacewatch observations from Oct. 14) and the following orbital elements appear on MPEC 2007-V75.

T = 2007 July 31.744 TT  ω = 277.^{o}971
e = 0.32388  Ω = 100.245 }2000.0
q = 2.74859 AU  i = 2.449
a = 4.06522 AU n^{o} = 0.120248 P = 8.2 years
2007 VA_{85}

An apparently asteroidal object of mag 18.5–19 discovered by the LIN-EAR survey on Nov. 4.1 UT has the shortest known period (P = 7.6 yr) for a sun-orbiting natural object with a retrograde orbit (i = 132^{o}.6). Attempts to find cometary activity have been unsuccessful. The available astrometry and orbital elements appear on MPEC 2007-V73.

COMET P/2007 T6 (CATALINA) Additional observations have show this comet (cf. IAUC 8881) to be

of short period, as indicated by the orbital elements on MPEC 2007-V56 (T = 2007 Aug. 19.586 TT, e = 0.50400, q = 2.23213 AU, ω = 335^{o}.833, Ω = 102^{o}.636, i = 22^{o}.148, equinox 2000.0, P = 9.55 yr).

2007 November 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green